20 eggs are in a refrigerator, 2 are bad, from the 4 are chosen at random to make a cake what is the probability that exactly one egg is bad



1
Expert's answer
2022-05-30T16:28:07-0400

Given

Total number of eggs in the refrigerator = 20

Number of bad eggs = 2

Number of good eggs = 20 - 2 = 18

"P(=1)=\frac {4!} {3!}0.1^10.9^{3}=0.29"